# Quantum mechanics provides the physical basis of teleological evolutions

### v4

@@ 2.2 First demonstration of the advanced knoledge rule
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem. Here we do not care about the fact that most NP problems are classically solved with a number of computation steps exponential in their size, so that a quadratic quantum speedup can be seen as a little advantage. Of course, for the present ends, we are just interested in having a quantum speedup with respect to what is classically possible.
@@
We shall see that there are further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
@@
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with their classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it.
@@
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm. We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of the quantum algorithm, all the information is selected by the initial measurement. Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
@@
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, that performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum algorithm. As we can see in Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum algorithm it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps. This is the physical correspondent of one of the logical causal loops whose mutually exclusive or is logically implied by the advanced knowledge rule.
@@
Eventually, we should take the quantum superposition of all the time-symmetrization instances. The time-symmetrized quantum description of the quantum algorithm (i.e. its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: So completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—it describes its physical counterpart.
@@ 2.4 Second demonstration of the advanced knowledge rule
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution. In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
@@
By the way, this result is reminiscent of Wheeler and Feynman’s [14] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on a reversibility of the quantum process that, unlike the classical one, is such in both an ideal and objective way.
2.5 Origin of the retrocausality
@@
A quantum algorithm is a relatively complex process. A natural question is whether the retrocausality implicit in the classical logic description of its overall result, i.e. the quantum speedup, derives from some more elementary part of it. Of course, in its ordinary quantum description, the unitary computation of the solution is an alternation of generations of quantum state superpositions and quantum interferences. Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality. Associating it with a causality that always goes in the same time-direction, as customary, must be consequent to the de facto no-retrocausality precept. Classical logic should also go along with the coexistence of causality and retrocausality.. Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
Let us now bring the above classical logic description of a state superposition to a quantum state superposition. Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
@@
The classical logic description of this can be as follows. In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question). Of course, this also provides the phase of the photon along the path followed. Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer. Then, it goes forward in time again taking the lower branch until reaching the detector in question and, this time, being detected. The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector. It should be clear that, also in this case, the involvement of retrocausality in the classical logic description is necessary.
@@
Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fundamental notion of quantum state superposition, the fact that the classical logic description of its generation necessarily involves retrocausality. By the way, note that this is not an idle speculation as it might even seem. Indeed, we are dealing with a form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
@@ 2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it. Note that we never needed to look inside the unitary evolution, for example $\hat{\mathbf{U}}_{1,2}$ of Tables I through IV of the Appendix. Thus, the need for this time-symmetrization does not only hold for quantum algorithms; there is the following generalization:
@@
Let us call it the quantum correlation time-symmetrization rule.
@@
The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup. However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables. Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says—see the Appendix. Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
@@
As shown in [5], the time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement. Note that this is exactly what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process equivalent in this case.
In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it. Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
@@
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
By the way, if the only possible hidden variables are retrocausal in character as the ones we are dealing with, this dismissal can be ascribed to the already mentioned de facto rejection of retrocausality and finalism (the form it has in philosophy, different from the present trivial form).
2.7 About the teleological character of the evolutions of the living
We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life. This section expands on some facts that, for compactness, were just mentioned, or omitted, in the Introduction.
Let us show that the possible teleological character of the evolution of the universe toward life must be inherited by the evolution of life on Earth. In principle, the latter evolution can be derived from that of the universe toward life by tracing out all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
@@ 2.7 About the teleological character of the evolutions of the living
In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would be the room for this) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
@@
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it will be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
@@
Let us begin with the evolution of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds. So, why anticipate something designed for flight many millions of years in advance?
@@
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually feathers turned out to be what was needed to conquer the air.
@@
All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. In our possibly biased opinion, this mutual causality explanation of the anticipation in question would be much more convincing than the customary causal one. See our preprint [6] for a more comprehensive discussion of the teleological notion.
@@
Another possible consequence is for the fact that, according to the present theory, a teleological evolution is an evolution endowed with quantum speedup. This could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [18]. Have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
Let us also recall what, here, is a very interesting observation of the French biologist Rémy Chauvin [19]. He pointed out the possibility that some strangely fast evolutionary changes of ants could happen before their being evolutionary advantages.
As is well known, this biologist merged important contributions to mainstream science with unconventional positions much criticized by colleagues. In addition to highlighting the possibility that some evolutionary changes are teleological in character, another unconventional position was his public interest in parapsychology (he wrote books on the subject). Note the parallel with the present situation. Let us consider the advance knowledge, on the part of the problem-solver, of any of the possible halves of the information about the solution of the problem she will read in the future. Besides being a completely technical notion at the basis of the present teleological notion (the problem-solver is a logical agent), under the quantum cosmological assumption it could well be the physical basis of the parapsychological notion of precognition.
@@
Of course, being under the quantum cosmological assumption, all the above is hypothetical. This is to decouple it from the teleological character of quantum algorithms, which we would consider certain.

### v7

@@ 2.2 First demonstration of the advanced knowledge rule
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
@@
Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem. We shall see that there are further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
By the way, let us underline the fact that the advanced knowledge rule tells us what can be the quantum speedup in the optimal quantum case. Of course, whether this quantum speedup is achieved, or achieved in part, or not at all, depends on the character of the unitary evolution that computes the solution. Therefore, we can say that the advanced knowledge rule tells us the room for the quantum speedup. Incidentally, achieving a quantum speedup may not be difficult: [15] shows that even a random quantum search for the solution yields just a suboptimal quantum speedup.
2.3 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
@@
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with their classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it; we shall also see that this time-symmetrization is mandatory to satisfy an atemporal symmetry implicit in this quantum description.
@@
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm.
Naturally, in the ordinary quantum description of the quantum algorithm, the initial measurement of the problem-setting selects the information that specifies both the problem-setting and the corresponding solution. To time-symmetrize this ordinary quantum description, we should evenly share the selection of this information between the initial measurement of the problem-setting and the final measurement of the solution. Since there is a plurality of possible ways of evenly sharing, we should take their quantum superposition.
@@
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. Time-symmetrically, the selection performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum description. As we can see in Table III of Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum description, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future (left of the bottom line of Table III diagram); she can then use this knowledge to compute the solution with fewer computation steps (bottom line of Table III diagram or Table IV diagram). The full Table III diagram is the physical correspondent of one of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
@@
Eventually, we should take the quantum superposition of all the time-symmetrization instances. The time-symmetrized quantum description of the quantum algorithm (i.e. its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: Thus completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—more precisely it describes its physical counterpart.
@@ 2.4 Second demonstration of the advanced knowledge rule
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it becomes simultaneous with the final measurement of the solution. At this point, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
@@
By the way, this result is reminiscent of Wheeler and Feynman’s [16] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on the reversibility of unitary quantum evolutions, which, unlike classical reversibility, is such in both an ideal and objective way.
2.5 Origin of the mutual causality
@@
A quantum algorithm is a relatively complex process. A natural question is whether the mutual causality implicit in the classical logic description of its overall result, i.e. the quantum speedup, derives from some more elementary part of it. Of course, in its ordinary quantum description, the unitary computation of the solution is an alternation of generations of quantum state superpositions and quantum interferences. Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
Let us show that the classical logic description of the unitary generation of a quantum state superposition necessarily involves retrocausality.
@@
Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
@@
The classical logic description of this quantum superposition can be as follows. In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question). Of course, this also provides the phase of the photon along the path followed. Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer. Then, it goes forward in time again taking the lower branch until reaching the detector in question, and this time, be detected. The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector. It should be clear that the classical logic description of this unitary generation of a quantum state superposition necessarily involves the coexistence of the two time directions of causality along the same unitary process.
@@
Summarizing, at the root of the form of mutual-causality we are dealing with there would be the fundamental notion of quantum state superposition. By the way, note that this is not an idle speculation as it might even seem. Indeed, we are dealing with a form of coexistence of the two time-directions of causality with huge consequences, such as providing a physical basis for the notion of teleological evolution.
@@ 2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have seen that the ordinary quantum description of a quantum algorithm whose solution is an invertible function of the problem-setting and is determined with certainty is incomplete and must be completed by time symmetrizing it. One can see that the necessity of this time-symmetrization does not depend on the fact that we are dealing with a quantum algorithm; there is this immediate generalization:
@@
Indeed, as one can see, to this kind of unitary evolution one can apply the argument for which it must be time-symmetrized to satisfy the atemporal symmetry implicit in it. Let us call this rule the quantum correlation time-symmetrization rule.
@@
In the case that the unitary evolution builds up one-to-one correlation between two initially unrelated observables, as in quantum search algorithms, this building up can always occur with quadratic quantum speedup, in the sense that there is always the room for its occurring with quantum speedup. This is what the time-symmetrization of its ordinary quantum description says—see Section (3) of the Appendix. Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
@@
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what should preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality with the two measurement apparatuses aligned. Here, the unitary evolution spatially separates the subsystems of two already maximally entangled observables. With the two measurement apparatuses aligned and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
@@
As shown in [5], the time-symmetrization of the ordinary quantum description of this quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [7]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement. Note that this is exactly what Einstein et al required from their hidden variables. That was, of course, without resorting to any mutual-causality, but under a deterministic view of the measurement process equivalent in this case. Of course, this should also mean that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
@@
By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might also be destabilizing. Indeed, after highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [17] and Ekert’s quantum cryptography based on it [18], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables. In particular, the argument of the incompleteness of the quantum description has been forgotten.
2.7 The possible teleological character of the evolutions of the living
We have seen that the possibility of a teleological character of a quantum algorithm, of its occurring with quantum speedup and thus be teleological in character, depends only on the form of its ordinary quantum description: of the unitary evolution between the initial an the final measurement we need only to know the input and the output. We do not need to know its inside. All this defines the room for the quantum speedup, which is what is of interest here. The unitary evolution might even not be a quantum algorithm.
Let us show how the teleological character of quantum algorithms could be ported to the (macroscopic) evolutions of the living for which the teleological notion was originally conceived. The first thing one should observe is that quantum algorithms belong to the microscopic world of quantum phenomena and, as things are now, have durations of a few dozen of microseconds: more than that, there is quantum decoherence. The evolutions of life, instead, can have the size of the earth and have duration of billions of years. The distance between the two is immense but, it should be noted, it would be so under the customary classical vision of macroscopic reality. Instead, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, this distance vanishes. By the way, recall the perfect time-reversibility of unitary quantum evolutions.
Now there is a surprising fact. Under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm and, as such, it can occur with quantum speedup and thus be teleological in character, this in the sense that there is certainly the room for that.
Let us explain. That version of the principle is about the extraordinary theoretical fact that the least change of the value (for values) that the fundamental physical constants has in the universe we are in would have produced a relatively trivial universe unable to develop life. Under the quantum cosmological assumption, this can be represented as follows. We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life” ; the latter should assume, in particular, the values “absence of life” and “existence of sentient life”.
We should then consider a quantum superposition (or, indifferently, mixture) of all the possible universes, each with its own value of the fundamental constants, all obeying the same fundamental laws (but for the specific value of the fundamental constants). Of course, at the beginning of time, in this quantum superposition, these two observables must be unrelated: no matter the value of the fundamental constants, the value of life is “absence of life”. We should now consider the unitary evolution of this quantum superposition of all possible universes toward the value of life. At the time of today, this evolution must have built up maximal entanglement (say, one-to-one correlation) between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value. By the way, according to cosmology, at the time of today, the absence of life in the universes that cannot produce it would already be definitive.
We have seen that this kind of building up of quantum correlation can always occur with quantum speedup (in the sense that there is always the room for it) and therefore be teleological in character. In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; and, at the time of today, there are the sentient observers who measure the value of life, also selecting the value of the fundamental constants maximally entangled with it. By the way, since the latter value is never changed by the unitary evolution of the universe toward life, all is as if it was measured at the beginning of time. We could also think that, if the search for the value of life by the evolution of the universe is random, it would occur with quantum speedup.
@@ 2.7 The possible teleological character of the evolutions of the living
Summarizing, the evolution of the universe toward life could occur with quantum speedup and consequently be teleological in character. Let us show that this possibility would then be inherited by the evolution of life on Earth and the evolution of species.
In principle, the evolution of life on Earth can be derived from the evolution of the universe toward life by tracing out all the rest. Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former.
In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would always be the room for it) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
@@
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it would be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
@@
It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly. Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds. So, why anticipate something designed for flight many millions of years in advance?
@@
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually (by sheer coincidence) feathers turned out to be what was needed to conquer the air.
@@
All this (without the parentheses and its content) could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. In our possibly biased opinion, this mutual-causality explanation of the anticipation in question would be much more convincing than the customary causal explanation. Indeed, the feathers of the fossils of those non-flying dinosaurs are almost identical to those of their flying descendants. Our preprint [6] provides a more extensive discussion of the implications of the teleological notion.
@@
Another consequence of the present theory is that, according to it, a teleological evolution is an evolution endowed with quantum speedup. Of course, this is a feature of teleological evolutions completely absent in their historical notion. It could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [19]. As is well known, have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
@@
Of course, being under the quantum cosmological assumption, all the above is hypothetical. This is to decouple it from the teleological character of quantum algorithms, which we would consider watertight.
3 Discussion
First, let us show that the abandonment of the teleological perspective in the natural sciences was and is not without criticism.
An example is a famous sentence by the evolutionary biologist Haldane [23]:
The other is the position of the natural philosopher Nagel [24], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
It could be said that, in general, natural philosophy is still pro the teleological notion. The following is a quotation from the natural philosophers Allen and Neal [22]:
Teleological notions are largely considered ineliminable from modern biological sciences, such as evolutionary biology, genetics, medicine, ethology, and psychiatry, because they play an important explanatory role.
Against the above, there is the position of the physical science and its de facto rejection of the teleological notion. By the way, this is also the position officially adopted by the natural sciences.
In our possibly biased view, natural philosophy would respond better to empirical evidence and be less affected by what we would consider a de facto prejudice of science.